Grade 12th pass9 grade science

Write the equations of motion. What does each letter indicate?

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
DEAR STUDENT

v = u + at

s = ut + 1/2 at2

v2 = u2 + 2as

Where u – Initial velocity  v – final velocity  a – acceleration  t – time  s – Distance (Displacement)

THANKS